c) Consider John who consumes two goods, (X and Y), with prices 𝑃π‘₯ = 𝑁$35, 𝑃𝑦 = 𝑁$25 and income I =N$1500 i) Construct budget constraint [3 marks] ii) Draw John’s budget line with good X on the horizontal axis. [3 marks] iii) Use a graph to show the effect of an increase in income from N$1500 to N$2000. [3 marks] iv) What will happen to the slope of the budget line if the price of good X decreases to N$18? [5 marks]


1
Expert's answer
2021-05-10T14:44:41-0400

"P_x=N\$35"

"P_y=N\$25"

"Income=N\$1500"


(i)

Budget constraint

"XP_x+YP_y\le I"

"35X+25Y\le 1500"


(ii)


"35X+25Y\le 1500"

if X=0, Y=60

if Y=0,X=42.857


(iii)

If there is an increases in income the budget line shifts rightward from AB to A'B' and the new budget constraint is

"35X+25Y\le 2000"




if the price of good X decreases to N$18 then quantity demanded of good X will increase with no change in quantity demanded of good Y and the budget line tilts from CD to CEand the slope of the budget line falls.